63 divided by r = 9
please show the work.

then there
17 = 153 divided by q
I am stomped.

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

9=63/r
9r=63 multiply each side by r.
r=7 divide each side by 9.
.
Ed

Answer by stanbon(75887)

63 divided by r = 9
---
63/r = 9
Invert both sides to get:
r/63 = 1/9
Multiply both sides by 63 to get:
----
r = 63/9
r = 7
-------------------------
17 = 153 divided by q
---
17 = 153/q
Invert both sides to get:
q/153 = 1/17
Multiply both sides by 153 to get:
q = 153/17
q = 9
===============
Cheers,
Stan H.